# Adaptive Federated Learning with Auto-Tuned Clients

### v1

@@ A.1 Proof of Algorithm 1 under Assumption 1
We start from the smoothness of $f(\cdot)$, which we emphasize again that is obtained via Lemma 3, i.e., as a byproduct of our analysis, and we do not assume that $f_{i}$’s are globally $L$-smooth. Also note that Lemma 3 holds for $\text{conv}\{x_{0},x_{1},\dots\}$, i.e., the average of the local parameters visited by Algorithm 1.
@@
Taking expectation from the expression above, we get
@@
where in the equality, we added and subtracted $\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\nabla f(x_{t})$ in the inner product term, and also replaced $D_{t}=\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}$. Then, using (7) and (12)-(14), we have
@@
We now expand the terms in (17), and will bound each seprately, as follows.
@@
$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
@@
$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,k}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\,\mathbb{E}\Big{\|}\nabla f\Big{(}x_{t}\Big{)}\Big{\|}^{2}$ $\leqslant\mathcal{O}\Big{(}\frac{\Psi_{1}}{\sqrt{T}}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{2}\Psi_{2}}{T}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{3}\Psi_{2}}{\sqrt{T^{3}}}\Big{)}.$
@@ A.2 Proof of Algorithm 1 for convex case
Here, we provide an extension of the proof of (Malitsky and Mishchenko, 2020, Theorem 1) for distributed version. Note that the proof we provide here is not exactly the same version of $\Delta$-SGD presented in Algorithm 1; in particular, we assume no local iterations, i.e., $k=1$ for all $i\in[m]$, and every client participate, i.e, $p=1.$
We start with the extension of (Malitsky and Mishchenko, 2020, Lemma 1), which constructs a Lyapunov function for the distributed objective in (1).
@@
Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex and twice differentiable. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$, satisfy the following:
@@
Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex and twice differentiable. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$, satisfy the following:
$\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}\leqslant\frac{2DL^{2}}{2Lt+1},$ (38)
@@
$S_{t}^{i}$ $=\eta_{t}^{i}(1+\theta_{t}^{i})+\sum_{k=1}^{t-1}\alpha_{k}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}.$
Further, if $x^{\star}$ is any minimum of $f_{i}(x)$ for all $i\in[m]$, (38) implies convergence for problem (1).
@@
Observe that (39) is of the form $\alpha_{t}^{i}[f_{i}(x_{t}^{i})-f_{i}(x^{\star})]+\sum_{k=1}^{t-1}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]=\sum_{k=1}^{t}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]$. The sum of the coefficients equals:
@@
$\eta_{t}^{i}=\frac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f(x_{t}^{i})-\nabla f(x_{t-1}^{i})\|}\geqslant\frac{1}{2\tilde{L}_{i}}\geqslant\frac{1}{2L},\quad\forall i\in[m].$
where $\tilde{L}_{i}$ is defined in (10), which by definition is smaller than or equal to the global smoothness $L$. Also, notice that $\eta_{1}^{i}\theta_{1}^{i}=\frac{(\eta_{1}^{i})^{2}}{\eta_{0}^{i}}=(\eta_{1}^{i})^{2},$ assuming for simplicity that $\eta_{0}^{i}=1.$ Thus, going back to (43), we have
$S_{t}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}$ $\geqslant\sum_{k=1}^{t}\frac{1}{2\tilde{L}_{i}}+\frac{1}{4\tilde{L}_{i}^{2}}$
$=\frac{t}{2L_{i}}+\frac{1}{4L_{i}^{2}}\geqslant\frac{t}{2L}+\frac{1}{4L^{2}}.$
Finally, we have
@@
$\geqslant\frac{1}{m}\sum_{i=1}^{m}\Big{(}\frac{t}{2L}+\frac{1}{4L^{2}}\Big{)}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)},$
@@
$\frac{2DL^{2}}{2Lt+1}\geqslant\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}.$
A.3 Proof of adaptive gradient descent in nonconvex case: centralized
For the centralized version, we use the original step size from Malitsky and Mishchenko (2020) as in (4).
Theorem 6.
Running gradient descent with step size in (4) for nonconvex $f$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(x_{t})\|^{2}\leqslant\frac{8\tilde{L}}{3T}[f(x_{0})-f^{\star}],$
[…]
$\|\nabla f(x)-\nabla f(y)\|\leqslant\tilde{L}\|x-y\|\quad\forall x,y\in\mathcal{C}$
Proof.
Using Lemma 2, where with slight abuse of notation we interpret $f_{i}$ as $f$ and $\eta_{t}^{i}$ as $\eta_{t},$ we have
$f(x_{t+1})$ $\leqslant f(x_{t})+\langle\nabla f(x_{t}),x_{t+1}-x_{t}\rangle+\frac{1}{4\eta_{t}}\|x_{t+1}-x_{t}\|^{2}$
$=f(x_{t})+\langle\nabla f(x_{t}),-\eta_{t}\nabla f(x_{t})\rangle+\frac{1}{4\eta_{t}}\|x_{t+1}-x_{t}\|^{2}$
$=f(x_{t})-\eta_{t}\|\nabla f(x_{t})\|^{2}+\frac{\eta_{t}}{4}\|\nabla f(x_{t})\|^{2}$
$=f(x_{t})-\frac{3\eta_{t}}{4}\|\nabla f(x_{t})\|^{2}$
Rearranging, we have
$\|\nabla f(x_{t})\|^{2}\leqslant\frac{4}{3\eta_{t}}[f(x_{t})-f^{\star}+f^{\star}-f(x_{t-1})]$
Unfolding for $T$ iterations and using telescoping, we get
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(x_{t})\|^{2}\leqslant\frac{4}{3T\eta_{t}}[f(x_{0})-f^{\star}]\leqslant\frac{8\tilde{L}}{3T}[f(x_{0})-f^{\star}]$
as desired. ∎
@@ B.1 Ease of tuning
B.2 Illustration of the degree of non-iidness
As mentioned in the main text, the level of "non-iidness" is controlled using latent Dirichlet allocation (LDA) applied to the labels. This approach is based on Hsu et al. (2019) and involves assigning each client a multinomial distribution over the labels. This distribution determines the labels from which the client’s training examples are drawn. Specifically, the local training examples for each client are sampled from a categorical distribution over $N$ classes, parameterized by a vector $\mathbf{q}$. The vector $\mathbf{q}$ is drawn from a Dirichlet distribution with a concentration parameter $\alpha$ times a prior distribution vector $\mathbf{p}$ over $N$ classes.
To vary the level of "non-iidness," different values of $\alpha$ are considered: $0.01,0.1$, and $1$. A larger $\alpha$ corresponds to settings that are closer to i.i.d. scenarios, while smaller values introduce more diversity and non-iidness among the clients, as visualized in Figure 5. As can be seen, for $\alpha=1,$ each client (each row on $y$-axis) have fairly diverse classes, similarly to the i.i.d. case; as we decrease $\alpha,$ the number of classes that each client has access to decreases. With $\alpha=0.01,$ there are man y clients who only have access to a single or a couple classes.
B.3 Effect of different number of local iterations
In this subsection, we show the effect of different number of local epochs performed per client, while fixing the experimental setup to be the classification of CIFAR-100 datset, trained with a ResNet-50. We show for all three cases of the Dirichlet concentration parameter $\alpha\in\{0.01,0.1,1\}$. The result is shown in Figure 6.
As mentioned in the main text, instead of performing local iterations, we instead perform local epochs $E,$ similarly to Reddi et al. (2021); Li et al. (2020). All the results in Table 1 use $E=1,$ and in Figure 6, we show the results for $E\in\{1,2,3\}$. Note that in terms of $E$, the considered numbers might be small difference, but in terms of actual local gradient steps, they differ significantly.
As can be seen, using higher $E$ leads to slightly faster convergence in all cases of $\alpha.$ However, the speed up seems marginal compared to the amount of extra computation time it requires (i.e., using $E=2$ takes roughly twice more total wall-clock time than using $E=1$). Put differently, $\Delta$-SGD performs well with only using a single local epoch per client ($E=1$), and still can achieve great performance.
Figure 6:
B.4 Additional experiments using different number of local data per client
Then, in the server aggregation step (line 13 of Algorithm 1), we compute the weighted average of the local parameters, instead of the plain average, as suggested in McMahan et al. (2017), and used for instance in Reddi et al. (2021).
We experiment in two settings: CIFAR-10 dataset classification trained with a ResNet-18 He et al. (2016), with Dirichlet concentration parameter $\alpha=0.1,$ which was the setting where the algorithms were fined-tuned using grid search; please see details in Section 4. We also provide results for CIFAR-100 dataset classification trained with a ResNet-50, with the same $\alpha.$
@@ B.4 Additional experiments using different number of local data per client
$\alpha=1$ SGD 80.7${}_{\downarrow(0.0)}$ 53.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.0)}}$
SGD ($\downarrow$)↓ ) 78.8${}_{\downarrow(1.9)}$ 53.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.9)}}$
SGDM ${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.7)}}$ 53.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.6)}}$
SGDM ($\downarrow$)↓ ) 66.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.1)}}$ 53.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.4)}}$
Adam 79.9${}_{\downarrow(0.8)}$ 51.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.4)}}$
Adagrad 79.3${}_{\downarrow(1.4)}$ 44.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.0)}}$
SPS 64.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.3)}}$ 37.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(20.3)}}$
$\Delta$-SGD 80.4${}_{\downarrow(0.3)}$ 57.5${}_{\downarrow(0.0)}$

### v3

@@ A.1 Proof of Algorithm 1 under Assumption 1
We also recall that we use ~$\tilde{\nabla}f_{i}(x)=\frac{1}{|\mathcal{B}|}\sum_{z\in\mathcal{B}}\nabla F_{i}(x,z)$ to denote the stochastic gradients with batch size $|\mathcal{B}|=b$. Note that whenever we use this shorthand notation in the theory, we imply that we draw a new independent batch of samples to compute the stochastic gradients.
We start from the smoothness of $f(\cdot)$, which we note again is obtained via Lemma 3, i.e., as a byproduct of our analysis, and we do not assume that $f_{i}$’s are globally $L$-smooth. Also note that Lemma 3 holds for $\{x_{0},x_{1},\dots\}$, i.e., the average of the local parameters visited by Algorithm 1.
@@
Taking expectation from the expression above, where we overload the notation such that the expectation is both with respect to the client sampling randomness (indexed with $i$) as well as the data sampling randomness (indexed with $z$), that is $\mathbb{E}(\cdot):=\mathbb{E}_{\mathcal{F}_{t}}\left[\mathbb{E}_{i}\left[\mathbb{E}_{z}[\cdot\mid\mathcal{F}_{t},i]|\mathcal{F}_{t}\right]\right],$ where $\mathcal{F}_{t}$ is the natural filtration, we have:
@@
where in the equality, we added and subtracted $\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}\nabla f(x_{t})$ in the inner product term, and also replaced $D_{t}:=\frac{1}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\eta_{t,k}^{i}$. Then, using (7) and (13)-(15), we have
@@
We now expand the terms in (18) and bound each term separately, as follows.
@@
$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{2K\tilde{L}^{2}}{D_{t}m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
@@
$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{[}\mathbb{E}\Big{\|}\tilde{\nabla}f_{i}(x_{t,\ell}^{i})-\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}+\mathbb{E}\Big{\|}\nabla f_{i}(x_{t,\ell}^{i})\Big{\|}^{2}\Big{]}$
$\leqslant\frac{4K\tilde{L}^{3}}{m}\sum_{i=1}^{m}\sum_{k=0}^{K-1}k(\eta_{t,k}^{i})^{2}\sum_{\ell=0}^{k-1}(\eta_{t,l}^{i})^{2}\Big{(}\frac{\sigma^{2}}{b}+G^{2}\Big{)}$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\,\mathbb{E}\big{\|}\nabla f(x_{t})\big{\|}^{2}$ $\leqslant\mathcal{O}\Big{(}\frac{\Psi_{1}}{\sqrt{T}}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{2}\Psi_{2}}{T}\Big{)}+\mathcal{O}\Big{(}\frac{\tilde{L}^{3}\Psi_{2}}{\sqrt{T^{3}}}\Big{)}.$
@@ A.2 Proof of Algorithm 1 for convex case
Here, we provide an extension of the proof of Malitsky & Mishchenko (2020, Theorem 1) for the distributed version. Note that the proof we provide here is not exactly the same version of $\Delta$-SGD presented in Algorithm 1; in particular, we assume no local iterations, i.e., $k=1$ for all $i\in[m]$, and every client participate, i.e, $p=1$, without stochastic gradients.
We start by constructing a Lyapunov function for the distributed objective in (1).
@@
Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex with locally Lipschitz gradients. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$ with full batch, satisfy the following:
@@
Let $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex with locally Lipschitz gradients. Then, the sequence $\{x_{t}\}$ generated by Algorithm 1, assuming $k=1$ and $p=1$ with full batch, satisfy the following:
$\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}\leqslant\frac{2D\hat{L}^{2}}{2\hat{L}t+1},$ (39)
@@
$S_{t}^{i}$ $=\eta_{t}^{i}(1+\theta_{t}^{i})+\sum_{k=1}^{t-1}\alpha_{k}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i},$
and $\hat{L}$ is a constant. Further, if $x^{\star}$ is any minimum of $f_{i}(x)$ for all $i\in[m]$, (39) implies convergence for problem (1).
@@
Observe that (40) is nonnegative by the definition of $\eta_{t}^{i},$ implying the iterates $\{x_{t}^{i}\}$ are bounded. Now, define the set $\mathcal{S}:=\mathrm{conv}(\{x_{0}^{i},x_{1}^{i},\dots\})$, which is bounded as the convex hull of bounded points. Therefore, $\nabla f_{i}$ is Lipschitz continuous on $\mathcal{S}$. That is, there exist $\hat{L}_{i}$ such that
$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leqslant\hat{L}_{i}\|x-y\|,\quad\forall x,y\in\mathcal{S}.$
Also note that (40) is of the form $\alpha_{t}^{i}[f_{i}(x_{t}^{i})-f_{i}(x^{\star})]+\sum_{k=1}^{t-1}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]=\sum_{k=1}^{t}\alpha_{k}^{i}[f_{i}(x_{k}^{i})-f_{i}(x^{\star})]$. The sum of the coefficients equals:
@@
$\eta_{t}^{i}=\frac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f(x_{t}^{i})-\nabla f(x_{t-1}^{i})\|}\geqslant\frac{1}{2\hat{L}_{i}},\quad\forall i\in[m].$
Also, notice that $\eta_{1}^{i}\theta_{1}^{i}=\frac{(\eta_{1}^{i})^{2}}{\eta_{0}^{i}}=(\eta_{1}^{i})^{2},$ assuming for simplicity that $\eta_{0}^{i}=1.$ Thus, going back to (44), we have
$S_{t}^{i}=\sum_{k=1}^{t}\eta_{k}^{i}+\eta_{1}^{i}\theta_{1}^{i}$ $\geqslant\sum_{k=1}^{t}\frac{1}{2\hat{L}_{i}}+\frac{1}{4\hat{L}_{i}^{2}}$
$=\frac{t}{2\hat{L}_{i}}+\frac{1}{4\hat{L}_{i}^{2}}\geqslant\frac{t}{2\hat{L}}+\frac{1}{4\hat{L}^{2}},$
where $\hat{L}=\max_{i}\hat{L}_{i}.$ Finally, we have
@@
$\geqslant\frac{1}{m}\sum_{i=1}^{m}\Big{(}\frac{t}{2\hat{L}}+\frac{1}{4\hat{L}^{2}}\Big{)}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)},$
@@
$\frac{2D\hat{L}^{2}}{2\hat{L}t+1}\geqslant\frac{1}{m}\sum_{i=1}^{m}\big{(}f_{i}(\tilde{x}_{t}^{i})-f_{i}(x^{\star})\big{)}.$
@@ B.1 Ease of tuning
B.2 Effect of different number of local iterations
In this subsection, we show the effect of the different number of local epochs, while fixing the experimental setup to be the classification of CIFAR-100 dataset, trained with a ResNet-50. We show for all three cases of the Dirichlet concentration parameter $\alpha\in\{0.01,0.1,1\}$. The result is shown in Figure 5.
As mentioned in the main text, instead of performing local iterations, we instead perform local epochs $E,$ similarly to (Reddi et al., 2021; Li et al., 2020). All the results in Table 1 use $E=1,$ and in Figure 5, we show the results for $E\in\{1,2,3\}$. Note that in terms of $E$, the considered numbers might be small difference, but in terms of actual local gradient steps, they differ significantly.
As can be seen, using higher $E$ leads to slightly faster convergence in all cases of $\alpha.$ However, the speed-up seems marginal compared to the amount of extra computation time it requires (i.e., using $E=2$ takes roughly twice more total wall-clock time than using $E=1$). Put differently, $\Delta$-SGD performs well with only using a single local epoch per client ($E=1$), and still can achieve great performance.
Figure 5: Effect of the different number of local epochs.
B.3 Additional experiments using different number of local data per client
@@
$\alpha=0.1$ SGD 80.7↓(0.0) 53.5↓(4.0)
SGD ($\downarrow$)↓ ) 78.8↓(1.9) 53.6↓(3.9)
SGDM 75.0↓(5.7) 53.9↓(3.6)
SGDM ($\downarrow$)↓ ) 66.6↓(14.1) 53.1↓(4.4)
Adam 79.9↓(0.8) 51.1↓(6.4)
Adagrad 79.3↓(1.4) 44.5↓(13.0)
SPS 64.4↓(16.3) 37.2↓(20.3)
$\Delta$-SGD 80.4↓(0.3) 57.5↓(0.0)